OpenStudy (anonymous):

What was the effects of the recent Fukushima nuclear leak in Japan? @whpalmer4

OpenStudy (whpalmer4):

3 ruined nuclear reactors, enormous economic costs, extensive (and in some cases ongoing) releases of radioactive I-131, Cs-134 and Cs-137 into air, ground and seawater, forced evacuation of hundreds of thousands of people, a sizable area around the plant still cordoned off, shutdown of all but 2 nuclear power reactors in Japan, etc.

OpenStudy (anonymous):

Thanks xx

OpenStudy (anonymous):

@whpalmer4 how long the effect of raiocative leaks will affect the environment . I need a half life diagram

OpenStudy (whpalmer4):

http://www.wolframalpha.com/input/?i=half+life+iodine-131%2C+cesium-134%2C+cesium-137 The time for radioactive decay of an element to reach a given percentage of the original can be determined from this formula: \[\% = e^{-\lambda t}\]where \[\lambda = \frac{0.693}{t_{1/2}}\]and \(t_{1/2}\) is the half-life. So, if you want to find the time it takes for a release of I-131 to decay to 10% of the original value: \[10\% = 10/100 = 0.1 = e^{-0.693t/8.0252} \]Take natural log of each side \[\ln 0.1 = \frac{-0.693t}{8.0252}\]\[\frac{8.0252*\ln 0.1}{-0.693} = t = 26.7 \text{ days}\] However, some of the leaks are continuing, and no one can say with any accuracy how long that will continue.

OpenStudy (anonymous):

i still don't know how i'm going to do the graph @whpalmer4

OpenStudy (whpalmer4):

I don't know what your diagram is supposed to look like. Perhaps if you were more explicit about what you expect a half-life diagram to be, I could tell you how to construct it.

OpenStudy (anonymous):

it has to be a graph

OpenStudy (whpalmer4):

A graph of what, precisely?

OpenStudy (anonymous):

it's a graph with a line

OpenStudy (whpalmer4):

A graph showing what relationship? What are the labels on the axes? What data is being plotted?

OpenStudy (anonymous):

Data beinf plotted

OpenStudy (whpalmer4):

Sorry, I don't know how you're going to the graph either, if you can't describe what you are trying to graph. Good luck!